OpenStudy (anonymous):

A tree 7 ft tall grows an average of 6 in. each year. Which equation models the tree’s height h after x years?

OpenStudy (alekos):

Do you have any ideas?

OpenStudy (anonymous):

h= -0.5 x + 7 h=0.5x +7 h=-6x +7 h=6x +7

OpenStudy (alekos):

are they you're choices?